1. In 1000 words, describe the Hardy- Weinberg Equilibrium. What are the assumptions and predictions? What is the relationship of the H-W equilibrium to population genetics & evolution?
Answer #1

Answere :

An English mathmematician G.H. Hardy and a German physician W. Weinberg, in the year 1908, independently developed a method to relate evolutionary changes and population genetics, popularly known as Hardy-Weinberg equilibrium law. This law states that the relative frequency of alleles in the population of sexually reproducing organisms remains constant from generation to generation as long as :

(i) the population is large enough and changes in allele frequencies due to chance are insignificant. It should be closed i.e no immigration and no emigration occur.

(ii) random mating occurs.

(iii) mutation do not occur.

(iv) No natural selection operates.

Hardy and Weinberg, through this law, have been able to explain why a population as a whole with all its variation and genotypic frequencies continue to remain unchanged for many generations. Mendelian genetic variation, in Hardy-Weinberg equation, are summed up as -

Assuming that a gene has two alleles A and a and if p is the frequency of occurence of dominant allele A, and q is the frequency of occurrence of recessive allel a in the parental generation then, the genotypic frequencies expected in the offspring as per Hardy-Weinberg equation are represented as :

(p+q) = p2 + 2pq + q2 =1 = gene frequency of the total population.

where, p2 (or p square) = frequency of occurence of individuals with homozygous dominant alleles (AA)

  2pq = frequency of occurence of heterozygous individuals (Aa)

  q2 (or q square) = frequency of occurence of individuals with homozygous recessive alleles (aa).